# A Chinese restaurant process for multiset permutations

### v2

@@ 4 The distribution of the number of cycles
An example where Theorem 4 holds is obtained by setting $\theta=1$, $n_{t}=\lfloor Ct^{\alpha}\rfloor$, for constants $C>0$, $\alpha>0$, where $\lfloor x\rfloor$ is the smallest integer greater or equal to $x$. Then, $n_{t}=Ct^{\alpha}+O(1)$ and $N_{t}=Ct^{\alpha+1}/(\alpha+1)+O(t^{\alpha})$ and the assumptions of Theorem 4 hold. The expectation and variance are asymptotically
${\mathbb{E}}(K_{t})\sim(\alpha+1)\log t\quad{\rm Var}(K_{t})\sim(\alpha+1)\log t,$
@@ 5 Discussion
see [2]. By considering the process of cycle counts $(C_{1}(n),C_{2}(n),\ldots,C_{n}(n))$, where $C_{i}(n)$ is the number of cycles of size $i$, we may write $K_{n}=\sum_{i=1}^{n}C_{i}(n)$, and, moreover, one may show that
$\frac{\sum_{i=1}^{\lfloor n^{t}\rfloor}C_{i}(n)-t\log n}{\sqrt{\log n}}$
is asymptotically standard Brownian motion; see [1]. If suitable approximations could be made to the component count process of multiset permutations, then similar progress might be made for the results obtained in this paper when $\theta=1$.

### v3

@@ 4 The distribution of the number of cycles
An example where Theorem 4 holds is obtained by setting $\theta=1$, $n_{t}=\lceil Ct^{\alpha}\rceil$, for constants $C>0$, $\alpha>0$, where $\lceil x\rceil$ is the smallest integer greater or equal to $x$. Then, $n_{t}=Ct^{\alpha}+O(1)$ and $N_{t}=Ct^{\alpha+1}/(\alpha+1)+O(t^{\alpha})$ and the assumptions of Theorem 4 hold. The expectation and variance are asymptotically
${\mathbb{E}}(K_{t})\sim(\alpha+1)\log t,\quad{\rm Var}(K_{t})\sim(\alpha+1)\log t.$
Another example for $\theta=1$ is obtained through regular variation [BGT]. A positive, measurable function $\ell(x)$, defined on $\eta,\infty)$ for $\eta$ real, is slowly varying if
$\lim_{x\to\infty}\frac{\ell(\lambda x)}{\ell(x)}=1,\quad\forall\lambda>0.$ (32)
The Uniform Convergence Theorem [BGT] provides that the convergence in (32) is uniform on compact $\lambda$-sets in $(0,\infty)$. A positive, measurable function defined on $\eta,\infty)$, is regularly varying with real index $\alpha$ if
$\lim_{x\to\infty}\frac{f(\lambda x)}{f(x)}=\lambda^{\alpha},\quad\forall\lambda>0.$
A regularly varying function with index $\alpha$ can be written as
$f(x)=x^{\alpha}\ell(x),$
where $\ell(x)$ is a slowly varying function. Suppose that $\ell(x)$ is locally bounded in $\eta,\infty)$, i.e. bounded on compact subsets of $\eta,\infty)$, and $\alpha>-1$. Karamata’s theorem gives
$\int_{\eta}^{t}x^{\alpha}\ell(x)\,dx\sim t^{\alpha+1}\ell(t)/(\alpha+1),\quad t\to\infty.$
We further suppose that $\eta\leq 1$, $\alpha\geq 0$, with $\lim_{x\to\infty}\ell(x)=\infty$ in the case $\alpha=0$, and define
$n_{t}=\Big\lceil\int_{t}^{t+1}x^{\alpha}\ell(x)\,dx\Big\rceil,\quad t\geq 1.$
The Uniform Convergence Theorem implies that
$\sup_{t\leq x\leq t+1}\left|\frac{\ell(x)}{\ell(t)}-1\right|\leq\sup_{1\leq\lambda\leq 2}\left|\frac{\ell(\lambda t)}{\ell(t)}-1\right|=o(1),$
from which
$\int_{t}^{t+1}x^{\alpha}\ell(x)\,dx$ $=$ $(1+o(1))\ell(t)\int_{t}^{t+1}x^{\alpha}\,dx$
$=$ $(1+o(1))\ell(t)(t^{\alpha}+o(t^{\alpha}))$
and
$n_{t}\sim t^{\alpha}\ell(t).$
Moreover, Karamata’s theorem gives us
$N_{t-1}\sim t^{\alpha+1}\ell(t)/(\alpha+1).$
Hence, the conditions of Theorem 4 hold.
We will now show that $n_{t}$ may be chosen in such a way that the conclusion of the central limit theorem does not hold for $K_{t}$. Suppose that $\theta=1$ and the $n_{t}$ are chosen such that
$N_{t-1}=o(n_{t}).$ (33)
By (24) and (25), we have
${\mathbb{E}}(X_{t})\sim\frac{n_{t}}{N_{t-1}},\quad{\rm Var}(X_{t})\sim\left(\frac{n_{t}}{N_{t-1}}\right)^{2}.$
We further choose the $n_{t}$ such that
$\sum_{s=1}^{t-1}{\rm Var}(X_{s})=o\left({\rm Var}(X_{t})\right),$ (34)
A particular example of such a sequence is
$n_{t}=\lceil e^{t^{3}}\rceil=e^{t^{3}}+O(1),$
from which
$N_{t}$ $=$ $\sum_{s=1}^{t}e^{s^{3}}+O(t)$
$=$ $e^{t^{3}}\left(1+\sum_{s=1}^{t-1}e^{s^{3}-t^{3}}\right)+O(t)$
$=$ $e^{t^{3}}\left(1+O\left(te^{(t-1)^{3}-t^{3}}\right)\right)+O(t)$
$=$ $e^{t^{3}}(1+o(1)).$
Clearly, (33) holds and we also have
$\sum_{s=1}^{t-1}{\rm Var}(X_{s})$ $=$ $O\left(\sum_{s=1}^{t-1}e^{2s^{3}-2(s-1)^{3}}\right)$
$=$ $O\left(\sum_{s=1}^{t-1}e^{6s^{2}-6s}\right)$
$=$ $O\left(te^{6(t-1)^{2}-6(t-1)}\right)$
$=$ $o\left(e^{6t^{2}-6t}\right)$
and so (34) holds, as well, because ${\rm Var}(X_{t})\sim e^{6t^{2}-6t+2}$.
Assuming (33) and (34), we have
$\frac{K_{t}-{\mathbb{E}}(K_{t})}{\sqrt{{\rm Var}(K_{t})}}=\frac{X_{t}-{\mathbb{E}}(X_{t})}{\sqrt{{\rm Var}(X_{t})}}(1+o(1))+\frac{\sum_{s=1}^{t-1}(X_{s}-{\mathbb{E}}(X_{s}))}{\sqrt{{\rm Var}(K_{t})}}.$ (35)
Suppose that
$\frac{K_{t}-{\mathbb{E}}(K_{t})}{\sqrt{{\rm Var}(K_{t})}}\stackrel{{\scriptstyle d}}{{\to}}{\rm N}(0,1).$
By our assumptions, the variance of the last term of (35) is $o(1)$ and so by Slutsky’s theorem [durrett]
$\lim_{t\to\infty}{\mathbb{P}}\left(\frac{X_{t}-{\mathbb{E}}(X_{t})}{\sqrt{{\rm Var}(K_{t})}}\leq 0\right)=\lim_{t\to\infty}{\mathbb{P}}\left(\frac{K_{t}-{\mathbb{E}}(K_{t})}{\sqrt{{\rm Var}(X_{t})}}\leq 0\right)=1/2,$ (36)
We will show that (36) is impossible. The c.d.f. of $X_{t}$ is given by
$P(X_{t}\leq x_{t})$ $=$ $\sum_{k_{t}=0}^{x_{t}}{\mathbb{P}}(X_{t}=k_{t})$
$=$ $\frac{1}{\binom{N_{t}}{n_{t}}}\sum_{k_{t}=0}^{x_{t}}\binom{N_{t-1}-1+n_{t}-k_{t}}{n_{t}-k_{t}}$
$=$ $\frac{1}{\binom{N_{t}}{n_{t}}}\left(\sum_{j=0}^{n_{t}}\binom{N_{t-1}-1+j}{j}-\sum_{j=0}^{n_{t}-x_{t}-1}\binom{N_{t-1}-1+j}{j}\right)$
$=$ $\frac{1}{\binom{N_{t}}{n_{t}}}\left(\binom{N_{t}}{n_{t}}-\binom{N_{t}-x_{t}-1}{n_{t}-x_{t}-1}\right)$
$=$ $1-\frac{(N_{t}-x_{t}-1)!/(n_{t}-x_{t}-1)!}{N_{t}!/n_{t}!}$
$\geq$ $1-\left(\frac{N_{t}-x_{t}-1}{N_{t}}\right)^{N_{t-1}},$ (39)
where we have used (22) at (4), (17) at (4), and $(N_{t}-x_{t}-1-i)/(N_{t}-i)\leq(N_{t}-x_{t}-1)/N_{t}$ for all $0\leq i\leq N_{t-1}-1$ at (39). We therefore have
$\liminf_{t\to\infty}P(X_{t}\leq{\mathbb{E}}(X_{t}))$ $\geq$ $1-\lim_{t\to\infty}\left(\frac{N_{t}-{\mathbb{E}}(X_{t})-1}{N_{t}}\right)^{N_{t-1}}$
$=$ $1-\lim_{t\to\infty}\left(\frac{N_{t}-n_{t}(1+o(1)/N_{t-1}-1}{N_{t}}\right)^{N_{t-1}}$
$=$ $1-\lim_{t\to\infty}\left(1-(1+o(1))/N_{t-1}\right)^{N_{t-1}}$
$=$ $1-e^{-1}>1/2,$
contradicting (36).
@@ 5 Discussion
see [BH]. By considering the process of cycle counts $(C_{1}(n),C_{2}(n),\ldots,C_{n}(n))$, where $C_{i}(n)$ is the number of cycles of size $i$, we may write $K_{n}=\sum_{i=1}^{n}C_{i}(n)$. Moreover, [DP] shows that $B_{n}(\cdot)\stackrel{{\scriptstyle d}}{{\to}}B(\cdot)$, where
$B_{n}(t):=\frac{\sum_{i=1}^{\lfloor n^{t}\rfloor}C_{i}(n)-t\log n}{\sqrt{\log n}},\quad 0\leq t\leq 1,$
and $B(t)$ is standard Brownian motion. If suitable approximations could be made to the process of cycle counts for multiset permutations, then similar progress might be made for the results obtained in this paper when $\theta=1$.
{acknowledgement}
[title=Acknowledgments] The author thanks the anonymous referee and the Associated Editor for a careful reading of this paper and for helpful suggestions.